OpenStudy (anonymous):

Could somebody please help me study for a pre-algebra test? I have some questions.

OpenStudy (slaaibak):

We will be here to answer your questions :)

OpenStudy (anonymous):

Okay thanks! I really appreciate it.

OpenStudy (anonymous):

Okay. sooo, this is really broad, but I don't understand negative exponents.... At all.

OpenStudy (akshay_budhkar):

So you got some specific questions or do you want me or us to explain you negative exponents from scratch?

OpenStudy (anonymous):

Ummmmmmm. Why does a negative exponent make a fraction? I don't understand the basic concept.

OpenStudy (slaaibak):

Negative exponents simply means this: \[a^{-x} = {1 \over a^x}\]

OpenStudy (akshay_budhkar):

yea slaaibak got it

OpenStudy (anonymous):

Yes but why are they defined that way?

OpenStudy (akshay_budhkar):

it is simple

OpenStudy (akshay_budhkar):

do you know \[a^x \times a^y= a^{x+y}\]

OpenStudy (anonymous):

No.

OpenStudy (akshay_budhkar):

now say we want \[a^x \times a^y = 1\]

OpenStudy (anonymous):

Well yeah I do sorry.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

You'd made one negative!!! AHHH I GET IT!

OpenStudy (akshay_budhkar):

so when we solve that we get \[a^{x+y}= 1 = a^0\]

OpenStudy (akshay_budhkar):

thats good! You get it :D

OpenStudy (anonymous):

Teehee! Wow that just clicked. My math teacher didn't explain it that way.......

OpenStudy (akshay_budhkar):

its my pleasure helping you click it :D

OpenStudy (anonymous):

So another thing, how do you deal with #s with exponents in a fraction?

OpenStudy (slaaibak):

Give an example please?

OpenStudy (anonymous):

Ummmm. How do you make fractions with this equation thing?

OpenStudy (slaaibak):

type over

OpenStudy (anonymous):

...

OpenStudy (akshay_budhkar):

3 over 4 \[3 \over 4\]

OpenStudy (slaaibak):

Use this code: {1 over 2} would be \[{1 \over 2}\]

OpenStudy (anonymous):

... Um can I just say it?

OpenStudy (akshay_budhkar):

yea what is your doubt? i do not comprehend it

OpenStudy (slaaibak):

yeah sure

OpenStudy (anonymous):

Okay so if you had y^-5 times 2^3 over y^4 times x^2.

OpenStudy (anonymous):

Or something to that effect.

OpenStudy (akshay_budhkar):

\[ y^-5 \times 2^3 \over y^4 \times x^2.\]

OpenStudy (slaaibak):

\[{ y^{-5} * 2^3 \over y^4 * x^2}\]

OpenStudy (akshay_budhkar):

is that your query?

OpenStudy (anonymous):

Lol yeah.

OpenStudy (slaaibak):

You're faster than me xD

OpenStudy (akshay_budhkar):

i am not lol

OpenStudy (anonymous):

=.=

OpenStudy (anonymous):

?

OpenStudy (slaaibak):

If there's multiplication, you can take y^-5 and but it below the line, \[y^{-5} = {1 \over y^5}\] You can immediately write 2^3 as 8. So now you have: \[{1 \over y^5} * {8 \over x^2 * y^4}\]

OpenStudy (slaaibak):

put*, not but

OpenStudy (slaaibak):

Then, you can multiply similar variables with eachother and add the powers. it becomes: \[8 \over y^9 * x^2\]

OpenStudy (anonymous):

What are you talking about?

OpenStudy (akshay_budhkar):

about the question you asked

OpenStudy (slaaibak):

I made a spelling mistake in my post.

OpenStudy (anonymous):

Yeah but how did you get y^9?

OpenStudy (anonymous):

I know you have to add... Or subtract the exponents or something.

OpenStudy (slaaibak):

\[y^4 * y^5 = y^{5 + 4}\]

OpenStudy (anonymous):

OH okay.

OpenStudy (akshay_budhkar):

the main aim of such questions is to convert your negative exponents into positive by taking its reciprocal

OpenStudy (anonymous):

Um one more thing. How would you get {1 over 9} % of 90. Or something like that.

OpenStudy (anonymous):

Aw shoot. I messed up the thing.

OpenStudy (anonymous):

\[{1 \over 9}\]

OpenStudy (akshay_budhkar):

percent is always divided by 100

OpenStudy (anonymous):

Yes.

OpenStudy (akshay_budhkar):

so 1/9 percent is \[(1/9 ) \over 100 = 1\over900\]

OpenStudy (akshay_budhkar):

1/900

OpenStudy (anonymous):

Sure...

OpenStudy (akshay_budhkar):

of 90 is 1/900 of 90.. of implies multiplication

OpenStudy (akshay_budhkar):

so it is 1/900 times 90 = 90/900 = 1/10

OpenStudy (anonymous):

Yes yes I get it.

OpenStudy (anonymous):

That makes sense! Okay that's all! And I have to go. Thank you so much. ^.^

OpenStudy (akshay_budhkar):

it is my pleasure